OpenStudy (anonymous):

Solve the following system: 2^(x+3y) = 128 and 2^(3x–y) = 2

OpenStudy (mayankdevnani):

http://answers.yahoo.com/question/index?qid=20090316232049AA3GjhK you get your answer , plz go through this link @jahvoan

OpenStudy (hba):

Apply log.

OpenStudy (callisto):

128 = 2^7 So, 2^(x+3y) = 128 2^(x+3y) = 2^7 Since the bases on both sides are the same, you can equate the power x+3y = 7 <- This is equation 1 2^(3x–y) = 2 2^(3x–y) = 2^1 Since the bases on both sides are the same, you can equate the power 3x-y = 1 <- this is equation 2 You have two equations and two unknowns. You can solve them now.